# CO(1–0) imaging reveals 10-kiloparsec molecular gas reservoirs around star-forming galaxies at high redshift

### v1

@@ 3 Analysis and results
Fits to individual galaxies show that the point-source model (our null hypothesis) is preferred for 15 out of 19 sources. For four sources, exponential profiles with are preferred (see Table 2): these are AS2COS13.1, AS2COS14.1, AS2COS28.1555AS2COS28.1 has a companion offset by 20 kpc, which might contribute to the large apparent source size (Chen et al., 2022a)., and AS2UDS126.0. These four sources do not stand out from the rest of the sample in redshift, SFR or . We hypothesise that our detections of extended emission in these source are due to their lower-than-average redshifts compared to the whole sample (increasing the line signal), which also place the CO(1–0) line at frequencies where JVLA has particularly low noise. In particular, AS2COS13.1 and AS2COS14.1 have the highest S/N out of the entire sample.
@@ 3.1 Spectral stacking: no signatures for CO(1–0) inflows or outflows
Are the extended CO reservoirs associated with molecular outflows or inflows? Molecular outflows – indicated by high-velocity wings of the line profile - have been invoked to explain extended [C ii] reservoirs around high-redshift galaxies from the ALPINE sample (Ginolfi et al., 2020; Pizzati et al., 2020, 2023); while ALPINE galaxies have SFRs an order-of-magnitude lower than our sample (10–100 M⊙/yr), a similar mechanism might be at play at even more highly star-forming DSFGs.
To search for potential high-velocity components, we stack the spectra of all 19 CO(1–0)-detected galaxies. We adopt Cleaned cubes, made with natural weighting. As the half-light radius of the stacked emission is smaller than the VLA beam, we extract spectra from an aperture equal to the beam size. The stacked spectrum is well-fit by a single Gaussian profile with a line FWHM of 64070 km s-1 (Fig. 2). Although our data is continuum-subtracted, we still find a low positive excess ( 5 Jy, increasing towards higher frequencies) across the entire spectrum. This residual continuum emission is likely due to curvature of the spectrum: at the rest-frame 115 GHz, the combination of free-free and dust thermal emission can lead to a significant spectral curvature. The lack of obvious outflow features indicates that the bulk of CO(1–0) gas is close to the systemic velocity.
@@ 3.2 ALMA [CI](1–0) data
We supplement our CO(1–0) data with [C i](1–0) ALMA observations of seven galaxies from our sample which were recently detected in [C i](1–0) line by Frias Castillo et al. (2024). The [C i](1–0) line is considered an alternative tracer of molecular gas, including the “CO-dark” phase (e.g., at low metallicities). The Frias Castillo et al. (2024) ALMA observations had baselines extending up to 140 k, giving a slightly higher angular resolution (2.2-4.0 circularised beam FWHM) than our JVLA data.
@@
Stacking all the data, we find that the exponential profile with =0.280.10 (2.20.8 kpc) is preferred over the point-source, albeit at lower significance than for CO(1–0) (AIC2.5). The inferred mean [C i](1–0) size is 4025% smaller than the CO(1–0) one, although they are consistent within 1.7. We therefore put an upper limit on the [C i](1–0) size by taking the 95th percentile of the posterior ( 0.42 (3.3 kpc). We thus can not rule out the that [C i](1–0) and CO(1–0) have similar scale lengths.
@@ 3.3 PDR modelling
We now infer the thermodynamics of the molecular gas reservoirs using photo-dissociation region (PDR) modelling. Specifically, we estimate the gas density and far-UV irradiation using the PDRToolbox suite of uniform-density PDR models (Kaufman et al., 2006; Pound & Wolfire, 2008). As we will show in Section 4, the CO(1–0)/[C i](1–0) sizes inferred from the uv-plane stacking are consistent with the extended [C ii] and far-IR continuum emission seen in previous studies of DSFGs.
@@
What if CO(1–0) is suppressed by CMB? Assuming a 50% suppression (an extreme scenario at , e.g., da Cunha et al. 2013), the inferred far-UV illumination increases by 0.5 dex (blue dotted line).
What if the bulk of [C ii] is arising from ionised gas rather than from the PDRs? Lowering the PDR contribution to the [C ii] flux down to 20% causes the inferred far-UV illumination to increase by almost 1 dex (red dotted line).
What if [C i](1–0) is more compact than CO(1–0) (i.e., the ratio of atomic/molecular carbon increases at large radii)? We assume the [C i](1–0) scale length indicated by the tentative results in Section 3.2 and re-calculate the [C i](1–0) surface brightness at the CO(1–0) half-light radius accordingly. The inferred density increases to cm-3 (green dotted line).
@@
where is the CO(1–0) flux per solid angle predicted by PDRToolbox and is the luminosity distance. Marginalising over and then yields %.
@@ 4.1 Sizes of extended molecular gas reservoirs in high-redshift DSFGs
Figure 5 places our measurement of CO(1–0) sizes in context of other high-redshift observations. Specifically, we show the literature CO(1–0) (Ivison et al., 2011; Thomson et al., 2012; Dannerbauer et al., 2017; Frias Castillo et al., 2022; Stanley et al., 2023) and [C ii] measurements (Gullberg et al., 2018; Rybak et al., 2019; Mitsuhashi et al., 2021). We homogenise the data by adopting the empirical CO excitation ladder from Birkin et al. (2021) and ; for the [C ii] measurements, we assume =30 from Zanella et al. (2018).
Our inferred CO(1–0) size is comparable to CO(1–0) sizes reported in the literature for other highly star-forming galaxy populations, and the [C ii] reservoirs seen around high-redshift DSFGs. The true extent of the molecular gas reservoirs in DSFGs might be even larger than those inferred from our CO(1–0) stacking. First, gas becomes less metal-rich with increasing distance from the centre of the galaxy (e.g., Schimek et al. 2024), causing to increase towards the outskirts. Moreover, the elevated CMB temperature at high redshift can significantly suppress the CO(1–0) emission, particularly at low gas kinetic temperatures and densities (by up to 50%, e.g., da Cunha et al. 2013; Zhang et al. 2016). However, the CMB effect will be more limited if the molecular gas is mostly in the form of higher-density ISM in star-forming regions or satellite galaxies.
@@ 4.2 Molecular gas versus dust and stellar emission
Using high-resolution 870-m ALMA imaging of the SCUBA-2 sources in the UKIDSS/UDS field (from which our sample is partially drawn), Gullberg et al. (2019) found a half-light radius of 1.10.3 kpc, assuming an exponential profile666After applying a correction to account for a faint extended component (Smail et al., 2021).. Somewhat larger sizes (1.80.1 kpc) were inferred for the ALESS sample of DSFGs (Hodge et al., 2016, 2024), which span a comparable redshift range (median ) and 850-m flux ( mJy) to our galaxies. In either case, the 870-m continuum is a factor of a more compact than the CO(1–0) reservoir.
@@
The presence of an extended molecular gas dovetails with the extended low-surface brightness dust continuum reported by Gullberg et al. (2019). After stacking ALMA imaging of 153 DSFGs, Gullberg et al. (2019) reported that 13% of the 870-m continuum arises for an extended component with an assumed radius of 4 kpc, similar to our inferred CO(1–0) size. This extended continuum emission lends further support to the presence of significant metal-enriched gas and reservoirs around high-redshift DSFGs.
As the bulk of molecular gas is far outside the central star-forming region, can the high SFRs of DSFGs be sustained by smooth gas accretion from this reservoirs? We can estimate the required inward streaming velocity by dividing the half-light radius by the median gas depletion time ( Myr, Frias Castillo et al. 2023). We obtain km s-1, comparable to the line-of-sight velocity dispersions seen in high-resolution kinematic studies of DSFGs (10 - 100 km s-1, e.g., Hodge et al. 2012; Rybak et al. 2015; Swinbank et al. 2015; Rizzo et al. 2021, 2023; Birkin et al. 2024. The central star-formation can thus be supported by smooth streaming of gas within the galaxy without major mergers (see, e.g., Narayanan et al. 2015).
For the sizes of stellar emission, we adopt the 4.4-m half-light radius of kpc derived by Gillman et al. (2024) for the AS2COSMOS and AS2UDS DSFGs. A similar value (3.00.3 kpc) was found for the ALESS sample by Hodge et al. (2024). The CO(1–0) emission in our sample is a factor of 1.40.2 and 1.30.2 more extended than 4.4-m sizes from Gillman et al. (2024) and Hodge et al. (2024), respectively. This size discrepancy indicates that in DSFGs, the cold molecular gas appears to extend beyond the stellar emission of the galaxy.
@@
The discrepancy in the spatial extent of molecular gas versus star formation and stars might be even more pronounced than indicated from the CO(1–0) and 4.4-m imaging. First, molecular gas might be more extended than CO(1–0) emission, due to higher and increased CMB attenuation in the low-density outskirts of galaxies. Second, if the dust temperature decreases towards larger radii (e.g., Calistro Rivera et al. 2018) star-formation will be more centrally concentrated than the observed-frame 870-m continuum. Finally, stellar component is likely more compact than the near-infrared continuum; high obscuration in the central region will “inflate” the apparent near-infrared sizes compared to the actual stellar component (Sorba & Sawicki, 2018; Popping et al., 2022; Smail et al., 2023; Gillman et al., 2024).
@@ 4.3 Comparison to molecular gas in hydrodynamical simulations
For a comparison in Fig. 5, we take data from the SIMBA cosmological simulation (Davé et al., 2019). SIMBA successfully reproduces key parameters of high-redshift DSFGs (Lovell et al., 2021). We focus the snapshot from the default 100-cMpc box. We select galaxies with M⊙, i.e., comparable in mass to our sample. Specifically, we consider the molecular gas masses and the (total) gas half-light radii from the SIMBA catalogue. We note that SIMBA does not predict H2 directly; rather, the molecular gas is predicted by post-processing the simulation data (Davé et al., 2019). The gas sizes are determined by averaging two-dimensional projections along the three axes.
@@
4.4 The nature and thermodynamics of the extended CO(1–0) reservoirs
@@
The combination of high , and low indicates that molecular gas resides in compact, dense clouds, that are distributed sparsely on the outskirts of galaxies. Alternatively, the clouds might be concentrated in satellites with small angular separation from the host (although we note the lack of companions at 1 separation in JWST NIRCam imaging of DSFGs in Gillman et al. 2024; Hodge et al. 2024). In either scenario, the molecular gas on the galaxy outskirts is significantly clumpy rather than distributed in a smooth, diffuse “halo”.
@@ 5 Conclusions
We have presented results of stacking CO(1–0) observations of a sample of 19 high-redshift DSFGs. Comprising 80 hours in the most compact JVLA array configuration, this is the deepest such study to date. Our main findings are:
The stacked -plane data show an extended exponential profile with a half-light radius of 3.80.5 kpc. The inferred size is comparable to extended CO(1–0) and [C ii] reservoirs previously reported around high-redshift DSFGs.
@@
Applying the same procedure to ALMA [C i](1–0) observations, we find that [C i](1–0) appears marginally more compact (2.20.8 kpc) than CO(1–0), although both sizes are consistent within 2.
@@
The inferred sizes of CO(1–0) reservoir are significantly larger than the 870-m dust continuum, indicating that the bulk of molecular gas (50–80%) resides outside the FIR-bright star-forming region. CO(1–0) is somewhat larger than the 4.4-m stellar emission.
Ultimately, understanding the nature of extended molecular gas reservoirs around high-redshift galaxies will require a new class of large, single-dish telescopes, such as the proposed 50-metre Atacama Large Aperture Submillimeter Telescope (AtLAST) (Mroczkowski et al., 2024; Lee et al., 2024).
@@ Acknowledgements.
M.R. is supported by the NWO Veni project ”Under the lens” (VI.Veni.202.225). J.A.H. acknowledges support from the ERC Consolidator Grant 101088676 (VOYAJ). I.R.S. and A.M.S. acknowledge STFC grant ST/X001075/1. E.F.-J.A. acknowledges support from UNAM-PAPIIT project IA102023, and from CONAHCyT Ciencia de Frontera project ID: CF-2023-I-506. C.-C.C. acknowledges support from the National Science and Technology Council of Taiwan (111-2112M-001-045-MY3), as well as Academia Sinica through the Career Development Award (AS-CDA-112-M02). The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. This paper makes use of the following ALMA data: ADS/JAO.ALMA #2021.1.01324.S. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. The authors are thankful for the assistance from Allegro, the European ALMA Regional Center node in the Netherlands. We thank D. Narayanan, E. Pizzati and A. Schimek for sharing insights from theoretical models and simulations.References
@@
- Hodge et al. (2024) Hodge, J. A., da Cunha, E., Kendrew, S., et al. 2024, arXiv e-prints, arXiv:2407.15846
@@
- Mroczkowski et al. (2024) Mroczkowski, T., Gallardo, P. A., Timpe, M., et al. 2024, arXiv e-prints, arXiv:2402.18645
@@
Appendix A Inferred source sizes
Table 2 lists the source sizes inferred from the MCMC modelling, using 5-k radial bin size. For galaxies for which the extended source is not preferred by both AIC and BIC, we give upper limit on the source sized, based on the 95th percentile from the posterior of the exponential model.
Source ID | CO(1–0) | [C i](1–0) | |||
[arcsec] | [kpc] | [arcsec] | [kpc] | ||
AS2COS1.1 | 4.625 | 2.62 | 17 | 1.36 | 9 |
AS2COS8.1 | 3.581 | 1.05 | 8 | 0.75 | 6 |
AS2COS13.1 | 2.608 | 0.41 | 2.9 | — | — |
AS2COS14.1 | 2.921 | 0.74 | 5.8 | 0.98 | 7.7 |
AS2COS23.1 | 4.341 | 1.89 | 13 | 1.110.44 | 7.63.0 |
AS2COS28.1 | 3.097 | 0.66 | 5.1 | — | — |
AS2COS31.1 | 3.643 | 5.26 | 39 | — | — |
AS2COS54.1 | 3.174 | 1.60 | 12 | — | — |
AS2COS65.1 | 2.414 | 0.68 | 5.6 | 0.380.11 | 3.10.9 |
AS2COS66.1 | 3.247 | 2.09 | 17 | — | — |
AS2COS139.1 | 2.219 | 1.08 | 8 | — | — |
AS2UDS10.0 | 3.169 | 3.65 | 28 | — | — |
AS2UDS11.0 | 4.073 | 2.37 | 38 | — | — |
AS2UDS12.0 | 2.520 | 0.80 | 6.6 | 0.400.19 | 3.31.6 |
AS2UDS126.0 | 2.436 | 0.97 | 8.0 | 0.640.28 | 5.32.3 |
AEG2 | 3.668 | 1.43 | 12 | — | — |
CDFN1 | 3.149 | 0.80 | 6 | — | — |
CDFN2 | 4.422 | 1.35 | 9 | — | — |

### v2

@@ 3 Analysis and results
Fits to individual galaxies show that the point-source model (our null hypothesis) is preferred for 15 out of 19 sources. For four sources, exponential profiles with are preferred (see Table 1): these are AS2COS13.1, AS2COS14.1, AS2COS28.1444AS2COS28.1 has a companion offset by 20 kpc, which might contribute to the large apparent source size (Chen et al., 2022a)., and AS2UDS126.0. These four sources do not stand out from the rest of the sample in redshift, SFR or . We hypothesise that our detections of extended emission in these source are due to their lower-than-average redshifts compared to the whole sample (increasing the line signal), which also place the CO(1–0) line at frequencies where JVLA has particularly low noise. In particular, AS2COS13.1 and AS2COS14.1 have the highest S/N out of the entire sample.
@@ 3.1 Spectral stacking: no signatures for CO(1–0) inflows or outflows
Are the extended CO reservoirs associated with molecular outflows or inflows? Gas outflows – indicated by high-velocity wings of the line profile – have been invoked to explain extended [C ii] reservoirs around high-redshift galaxies from the ALPINE sample (Ginolfi et al., 2020; Pizzati et al., 2020, 2023). Although ALPINE galaxies have SFRs an order-of-magnitude lower than our sample (10–100 M⊙ yr-1), the same mechanism might be at play at even more highly star-forming DSFGs.
To search for potential high-velocity components, we stack the spectra of all 19 CO(1–0)-detected galaxies. We adopt Cleaned cubes, made with natural weighting. As the half-light radius of the stacked emission is smaller than the VLA beam, we extract spectra from an aperture equal to the beam size. The stacked spectrum is well-fit by a single Gaussian profile with a line FWHM of 59070 km s-1 (Fig. 2). Although our data is continuum-subtracted, we still find a low positive excess ( 5 Jy, increasing towards higher frequencies) across the entire spectrum. This residual continuum emission is likely due to curvature of the spectrum: at the rest-frame 115 GHz, the combination of free-free and dust thermal emission can lead to a significant spectral curvature. The lack of obvious outflow features indicates that the bulk of CO(1–0) is close to the systemic velocity.
@@ 3.2 ALMA [CI](1–0) data
We supplement our CO(1–0) data with [C i](1–0) ALMA observations of ten galaxies from our parent sample detected in [C i](1–0) line by Chen et al. (2022a, AS2COS31.1) and Frias Castillo et al. (2024, nine galaxies555The tenth galaxy detected in () – AS2COS66.1 – has only part of the [C i](1–0) line covered by the ALMA tuning; consequently, we do not include it in our analysis). The [C i](1–0) line is considered an alternative tracer of molecular gas, including the “CO-dark” phase (e.g., at low metallicities). The Frias Castillo et al. (2024) ALMA observations had baselines extending up to 140 k, giving a slightly higher angular resolution (2.2-4.0 circularised beam FWHM) than our JVLA data.
@@
Stacking all the data, we find that the exponential profile with =0.200.11 (1.60.8 kpc) is preferred over the point-source one, albeit at low significance (AIC2). The inferred mean [C i](1–0) size is 6025% smaller than the CO(1–0) one, as a 2.3 significance. Given the very weak preference for the exponential profile over the point-source one, we conservatively put an upper limit on the [C i](1–0) size by taking the 95th percentile of the posterior ( 0.33 (2.6 kpc). The [CI](1-0) therefore appears to be marginally less extended than CO(1-0).
@@ 3.3 PDR modelling
We now infer the thermodynamics of the molecular gas reservoirs using photo-dissociation region (PDR) modelling. Specifically, we estimate the gas density and far-UV irradiation using the PDRToolbox suite of uniform-density PDR models (Kaufman et al., 2006; Pound & Wolfire, 2008). As we will show in Section 4, the CO(1–0)/[C i](1–0) sizes inferred from the uv-plane stack are consistent with the extended component of the [C ii] and far-IR continuum emission seen in previous studies of DSFGs.
@@
What if the bulk of [C ii] is arising from ionised gas rather than from the PDRs? Lowering the PDR contribution to the [C ii] flux down to 20% causes the inferred far-UV illumination to increase by almost 1 dex (red dotted line in Fig. 4).
What if [C i](1–0) is more compact than CO(1–0) (i.e., the ratio of atomic/molecular carbon increases at large radii)? We assume the [C i](1–0) scale length indicated by the tentative results in Section 3.2 and re-calculate the [C i](1–0) surface brightness at the CO(1–0) half-light radius accordingly. The inferred density increases to cm-3 (green dotted line in Fig. 4).
@@
where is the CO(1–0) flux per solid angle predicted by PDRToolbox and is the luminosity distance. Marginalising over and then yields %. We discuss the context of this result in Section 4.4.
@@ 4 Discussion
In this Section, we compare the results of our stacking procedure and PDR modelling to other studies of high-redshift DSFGs (4.1), dust and stellar emission (4.2), and hydrodynamical simulations (4.3). Finally, we address the nature of extended CO(1–0) reservoirs (4.4).
@@ 4.1 Sizes of extended molecular gas reservoirs in high-redshift DSFGs
Figure 5 places our measurement of CO(1–0) sizes in context of other high-redshift observations. Specifically, we show the literature CO(1–0) (Ivison et al., 2011; Thomson et al., 2012; Dannerbauer et al., 2017; Frias Castillo et al., 2022; Stanley et al., 2023) and [C ii] measurements (Gullberg et al., 2018; Rybak et al., 2019; Mitsuhashi et al., 2021). We derive with (e.g., Calistro Rivera et al. 2018). For the [C ii]-based masses, we use , assuming =30 (Zanella et al., 2018).
Our inferred CO(1–0) size is comparable to CO(1–0) sizes reported in the literature for other highly star-forming galaxy populations, and the [C ii] reservoirs seen around high-redshift DSFGs. Our inferred sizes are also consistent with the CO(1–0) sizes reported by Stanley et al. (2023) for the (potentially) non-lensed sources from the Z-GAL sample (median 4.4 kpc). In contrast, our CO(1–0) size is larger than the typical extent of CO(3–2) emission in DSFGs—2.6 kpc—recently reported by Amvrosiadis et al. (2025). The true extent of the molecular gas reservoirs in DSFGs might be even larger than those inferred from our CO(1–0) stacking. First, gas becomes less metal-rich with increasing distance from the centre of the galaxy (e.g., Schimek et al. 2024), causing to increase towards the outskirts. Moreover, the elevated CMB temperature at high redshift can significantly suppress the CO(1–0) emission, particularly at low gas kinetic temperatures and densities (by up to 50%, e.g., da Cunha et al. 2013; Zhang et al. 2016). However, the CMB effect will be more limited if the molecular gas is mostly in the form of higher-density ISM in star-forming regions or satellite galaxies.
The CO(1–0) extent derived from our stacking is comparable to the sizes of [C ii] emission in high-redshift galaxies with similar SFR and molecular gas mass (Fig. 5, e.g., Gullberg et al. 2018; Rybak et al. 2019; Mitsuhashi et al. 2021), as well as those reported around star-forming galaxies from the CRISTAL sample ( kpc, Ikeda et al. 2025).
The extended [C ii] reservoirs around high-redshift galaxies have been hypothesised to arise primarily from atomic or ionised gas (e.g., Ikeda et al. 2025), perhaps deposited by outflows (e.g., Pizzati et al. 2020, 2023. However, the comparable extent of the CO(1–0) (which is associated only with molecular gas) indicates that molecular gas might be the dominant component of the extended gas halos.
@@ 4.2 Molecular gas versus dust and stellar emission
Using high-resolution 870-m ALMA imaging of the SCUBA-2 sources in the UKIDSS/UDS field (from which our sample is partially drawn), Gullberg et al. (2019) found a half-light radius of 1.10.3 kpc, assuming an exponential profile666After applying a correction to account for a faint extended component (Smail et al., 2021).. Somewhat larger sizes (1.80.1 kpc) were inferred for the ALESS sample of DSFGs (Hodge et al., 2016, 2025), which span a comparable redshift range (median ) and 850-m flux ( mJy) to our galaxies. In either case, the 870-m continuum is a factor of more compact than the CO(1–0) reservoir.
@@
The presence of an extended molecular gas dovetails with the extended low-surface brightness dust continuum reported by Gullberg et al. (2019), who after stacking ALMA imaging of 153 DSFGs, reported that 13% of the 870-m continuum arises from an extended component with an assumed radius of 4 kpc, which is similar to our measured CO(1–0) size. This extended continuum emission lends further support to the presence of significant metal-enriched gas and reservoirs around high-redshift DSFGs.
As the bulk of the molecular gas is far outside the central star-forming region, can the high SFRs of DSFGs be sustained by smooth gas accretion from this reservoirs? We can estimate the required inward streaming velocity by dividing the half-light radius by one half of the median gas depletion time ( Myr, Frias Castillo et al. 2023). We obtain km s-1, consistent with to the line-of-sight velocity dispersions seen in high-resolution kinematic studies of DSFGs (10 - 100 km s-1, e.g., Hodge et al. 2012; Rybak et al. 2015; Swinbank et al. 2015; Rizzo et al. 2021, 2023; Birkin et al. 2024; Amvrosiadis et al. 2025. The central star formation can thus be supported by smooth streaming of gas within the galaxy without major mergers (see, e.g., Narayanan et al. 2015).
For the sizes of stellar emission, we adopt the 4.4-m half-light radius of kpc derived by Gillman et al. (2024) for the AS2COSMOS and AS2UDS DSFGs. A similar value (3.00.3 kpc) was found for the ALESS sample by Hodge et al. (2025). The CO(1–0) emission in our sample is a factor of 1.40.2 and 1.30.2 more extended than the 4.4-m sizes from Gillman et al. (2024) and Hodge et al. (2025), respectively. This size discrepancy indicates that in DSFGs, the cold molecular gas appears to extend beyond the stellar emission of the galaxy.
@@
The discrepancy in the spatial extent of molecular gas versus star formation and stars might be even more pronounced than indicated from the CO(1–0) and 4.4-m imaging. First, molecular gas might be more extended than CO(1–0) emission, due to higher and increased CMB attenuation in the low-density outskirts of galaxies. Second, if the dust temperature decreases towards larger radii (e.g., Calistro Rivera et al. 2018) star formation will be more centrally concentrated than the observed-frame 870-m continuum. Finally, the stellar component is likely more compact than the near-infrared continuum; high obscuration in the central region will “inflate” the apparent near-infrared sizes compared to the actual stellar component (Sorba & Sawicki, 2018; Popping et al., 2022; Smail et al., 2023; Gillman et al., 2024).
@@ 4.3 Comparison to molecular gas in hydrodynamical simulations
For a comparison in Fig. 5, we take data from the SIMBA cosmological simulation (Davé et al., 2019). SIMBA successfully reproduces key parameters of high-redshift DSFGs (Lovell et al., 2021). We focus on the snapshot from the default 100-cMpc box. We select galaxies with M⊙, i.e., comparable in mass to our sample. Specifically, we consider the molecular gas masses and the (total) gas half-light radii from the SIMBA catalogue. We note that SIMBA does not predict H2 directly; rather, the molecular gas is predicted by post-processing the simulation data (Davé et al., 2019). The gas sizes are determined by averaging two-dimensional projections along the three axes.
@@
4.4 The nature of the extended CO(1–0) reservoirs
@@
The combination of high , and low indicates that molecular gas resides in compact, dense clouds, that are distributed sparsely on the outskirts of galaxies. Alternatively, the clouds might be concentrated in satellites with small angular separation from the host (although we note the lack of companions at 1 separation in JWST NIRCam imaging of DSFGs in Gillman et al. 2024; Hodge et al. 2025). In either scenario, the molecular gas on the galaxy outskirts would be significantly clumpy rather than distributed in a smooth, diffuse “halo”.
@@ 5 Conclusions
We have presented results of stacking CO(1–0) observations of a sample of 19 high-redshift DSFGs. Comprising 80 hours in the most compact JVLA array configuration, this is the deepest study of extended cold-gas reservoirs around early galaxies to date. Our main findings are:
The stacked -plane data of the CO(1–0) emission show an extended exponential profile with a half-light radius of 3.80.5 kpc. The inferred size is comparable to extended CO(1–0) and [C ii] reservoirs previously reported around high-redshift DSFGs.
@@
Applying the same procedure to ALMA [C i](1–0) observations, we find that [C i](1–0) appears more compact (2.6 kpc) than CO(1–0), although both sizes are consistent within 2. Five galaxies show extended [C i](1–0) emission, with half-light radii of 3.1–7.7 kpc.
@@
The inferred sizes of CO(1–0) cold-gas reservoir are significantly larger than the bulk of the 870-m dust continuum, but similar to the fainter, extended dust component that has been reported in these sources. The CO(1–0) is somewhat larger than the 4.4-m stellar emission. The bulk of the molecular gas (50–80%) resides outside the FIR-bright star-forming region and does not directly contribute to the observed star formation.
Our findings reveal that the bulk of molecular gas in high-redshift DSFGs resides in extended, 10-kpc scale reservoirs; only a small fraction of this gas directly contributes to the observed star formation.
@@ Acknowledgements.
M.R. is supported by the NWO Veni project ”Under the lens” (VI.Veni.202.225). J.A.H. acknowledges support from the ERC Consolidator Grant 101088676 (VOYAJ). I.R.S. and A.M.S. acknowledge STFC grant ST/X001075/1. E.F.-J.A. acknowledges support from UNAM-PAPIIT project IA102023, and from CONAHCyT Ciencia de Frontera project ID: CF-2023-I-506. C.-C.C. acknowledges support from the National Science and Technology Council of Taiwan (111-2112M-001-045-MY3), as well as Academia Sinica through the Career Development Award (AS-CDA-112-M02). The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. This paper makes use of the following ALMA data: ADS/JAO.ALMA #2021.1.01324.S. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. The authors are thankful for the assistance from Allegro, the European ALMA Regional Center node in the Netherlands. We thank D. Narayanan, E. Pizzati and A. Schimek for sharing insights from theoretical models and simulations. Finally, we thank the referee – P. Cox – for his detailed comments that helped us improve this manuscript.References
- Amvrosiadis et al. (2025) Amvrosiadis, A., Wardlow, J. L., Birkin, J. E., et al. 2025, MNRAS, 536, 3757
- Battisti et al. (2019) Battisti, A. J., da Cunha, E., Grasha, K., et al. 2019, ApJ, 882, 61
@@
- Hodge et al. (2025) Hodge, J. A., da Cunha, E., Kendrew, S., et al. 2025, ApJ, 978, 165